A ball m= 0.14 kg with initial velocity of 40 m/s is hit. it leaves the bat with a velocity of 60 m/s after 0.001 s. what is the impulse an average net force applied to the ball by the bat.


1
Expert's answer
2022-05-19T05:43:04-0400

The impulse

"\Delta p=m\Delta v=0.14*(60-40)=2.8\:\rm kg\cdot m/s"

The average force

"F=\frac{\Delta p}{\Delta t}=\frac{2.8}{0.001}=2800\:\rm N"